OpenStudy (minecraftgeek5):

The average hourly wage for part-time workers at a juice bar is $9.50. The actual hourly wage could differ from the average by as much as $1.50. How would I write this as an absolute value inequality

OpenStudy (minecraftgeek5):

I will medal and fan

OpenStudy (mathstudent55):

9.50 - 1.5 = 8 minimum salary 9.50 + 1.5 = 11 maximum salary \(8 \le x \le 11\) \(8- 9.5 \le x - 9.5 \le 11 - 9.5\) \(-1.5 \le x - 9.5 \le 1.5\) \(|x - 9.5| \le 1.5\)

OpenStudy (minecraftgeek5):

Thank you sooo much!!! This helped a lot ^_^

OpenStudy (mathstudent55):

You're welcome.